one pound of air at 15 psia, 300 f is heated at constant volume reversibly until its temperature becomes 600 F. Find P2,W,


1
Expert's answer
2021-12-14T15:08:52-0500

Given quantities:

"P_1=15 psia"

"T_1=300F = (300-32)*\frac{5}{9}+273=422K"

"T_1=600F = (600-32)*\frac{5}{9}+273=588K"


"\frac{P}{T}=const"


"\frac{P_1}{T_1}=\frac{P2}{T_2} \to P_2=\frac{T_2}{T_1}P_1=\frac{588}{422}*15=20.9 psia"